(a) What is the maximum torque on a 150-turn square loop of wire 18.0 cm on a side that carries a 50.0-A current in a 1.60-T field? _____ N-m

(b) What is the torque when θ is 10.9°? _____ N-m


Round your answers to the nearest 1 decimal place

1
Expert's answer
2021-05-28T07:18:50-0400

Explanations & Calculations


  • The Torque generated by a loop of a conductor placed in a uniform magnetic field is given by

"\qquad\qquad
\begin{aligned}
\small \tau&=\small BiNA\sin\theta
\end{aligned}"

  • The angle is measured with respect to the plane of the magnetic field such that at 90 degrees the torque achieves its maximum as "\small \sin 90 = 1".
  • That is when the loop positions itself perpendicular to the field.
  • Similarly, no torque is experienced when the loop is parallel to the field when "\small \theta = 0" as "\small \sin 0 =0".
  • Then,

"\qquad\qquad
\begin{aligned}
\small \tau_{max}&=\small BiNA\\
&= \small 1.60T\times50A\times150\times(0.18m)^2\\
&=\small \bold{388.8\,Nm}\\\\

\tau_{@\,\theta=10.9^0}&=\small 1.60\times50\times150\times(1.8)^2\times\sin10.9\\
&=\small \bold{73.5\,Nm}
\end{aligned}"